# Determination of the stably free cancellation property for orders

### v1

@@ Abstract.
Let $K$ be a number field, let $A$ be a finite-dimensional semisimple $K$-algebra, and let $\Lambda$ be an $\mathcal{O}_{K}$-order in $A$. We give practical algorithms that determine whether or not $\Lambda$ has the stably free cancellation property (SFC). As an application, we determine all finite groups $G$ of order at most $383$ such that the integral group ring $\mathbb{Z}[G]$ has SFC.
@@ 1. Introduction
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$. Let $A$ be a finite-dimensional semisimple $K$-algebra and let $\Lambda$ be an $\mathcal{O}$-order in $A$. For example, if $G$ is a finite group then the group ring $\mathcal{O}[G]$ is an order in the group algebra $K[G]$. For $k\geq 0$, let $\Lambda^{(k)}$ denote the (left) $\Lambda$-module that is the direct sum of $k$ copies of $\Lambda$. The order $\Lambda$ is said to have stably free cancellation (SFC) if every finitely generated stably free $\Lambda$-module is in fact free, that is, if $X$ is a finitely generated $\Lambda$-module such that $X\oplus\Lambda^{(n)}\cong\Lambda^{(n+m)}$ for some $m,n\geq 0$, then $X\cong\Lambda^{(m)}$.
@@
In this article, we show that the question of whether an arbitrary order $\Lambda$ has SFC or not can in principle always be decided algorithmically. In fact, we present three algorithms, all of which are practical in certain situations. The first of these, Algorithm 8.9, is only practical when $A$ is of ‘small’ dimension. The second, Algorithm 9.1, cannot prove that $\Lambda$ has SFC, but can be used to show that it fails SFC, and in this situation is often much faster than Algorithm 8.9. The third, Algorithm 10.3, uses fibre products to reduce the problem for the original order to that for an order in an algebra of smaller dimension as well as the computation of a certain subgroup of the unit group of a finite ring. These algorithms require several other algorithms of independent interest, including those that can be used to compute certain primary decompositions and unit groups of finite rings (see §11), and improved freeness tests for certain modules over orders (see §12).
Orders of particular interest are integral group rings $\mathbb{Z}[G]$ for $G$ a finite group. For instance, the property of whether or not $\mathbb{Z}[G]$ has SFC or LFC has applications in topology (see [Joh03], [Nic21b], [Nic20], for example) and in the determination of whether a finite Galois extension of $\mathbb{Q}$ has a normal integral basis or not (see [Cou94], [Cou98]).
@@
A special case of a result due to Swan and Fröhlich says that if $\mathbb{Z}[G]$ has SFC then $\mathbb{Z}[G/N]$ has SFC for every quotient $G/N$, or equivalently, if $\mathbb{Z}[G/N]$ fails SFC for some choice of $N$, then $\mathbb{Z}[G]$ also fails SFC (see Theorem 3.1 and Lemma 4.1). The following result is a special case of Theorem 9.4, in which the claim that each $\mathbb{Z}[G]$ fails SFC was proven using Algorithm 9.1; the claim that $\mathbb{Z}[G/N]$ has SFC for each proper quotient $G/N$ can be interpreted as saying that each $G$ is minimal with respect to quotients.
@@
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}$, or
@@
$\tilde{T}\times C_{2},\,\tilde{T}\times Q_{12},\,\tilde{T}\times Q_{20},\,Q_{8}\rtimes\tilde{T},\,G_{(96,66)},\,G_{(192,183)},\,G_{(480,962)}.$
@@
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}$,
@@
This represents a significant advance in the state-of-the-art, since before the results of this article, such a classification was only known for groups $G$ with $|G|\leq 31$ (for example, it was not known whether $\mathbb{Z}[G_{(32,14)}]$ had SFC or not). In fact, Theorem 1.4 is a special case of Theorem 13.2, which gives a classification for $|G|\leq 1023$, but with $27$ explicitly listed gaps. In other words, of the $11759892$ groups $G$ with $|G|\leq 1023$, there are now only $27$ for which it is not known whether $\mathbb{Z}[G]$ has SFC or not.
Nicholson [Nic24, Theorem 1.7] has now shown that $\mathbb{Z}[\tilde{T}\times C_{2}]$ in fact has LFC, using our result that it has SFC (i.e. part of Theorem 1.3) as an input. By combining this with further new results on integral group rings with LFC, and using Theorem 1.1 as a crucial input, he has given classifications of all integral group rings $\mathbb{Z}[G]$ with LFC and SFC, provided that either (i) $G$ has no quotient isomorphic to $\tilde{O}$ or $\tilde{I}$, or (ii) $G$ does have a quotient isomorphic to $C_{2}\times C_{2}$ (see [Nic24, Theorems A and B]). We remark that in the case of integral group rings, the LFC property discussed here is equivalent to the projective cancellation (PC) property discussed in loc. cit.
@@ Acknowledgements
The authors are grateful to John Nicholson for helpful comments and discussions, including suggestions for which integral group rings to investigate initially. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any author accepted manuscript version arising.
@@ 2.5. Locally free class groups
where $X_{1},X_{2},X_{3}\in g(\Lambda)$. This is now a well-defined commutative and associative operation, with identity element $[\Lambda]_{\mathrm{st}}$. Moreover, inverses exist since by (2.1), given $X\in g(\Lambda)$ there exists $X^{\prime}\in g(\Lambda)$ such that $X\oplus X^{\prime}\cong\Lambda^{(2)}$. Therefore $\mathrm{Cl}(\Lambda)$ is an abelian group, and is finite since $\mathrm{LF}_{1}(\Lambda)$ is finite.
@@ 3.1. Stably free and locally free cancellation
By (2.2), $\Lambda$ has SFC if and only if (3.1) holds for all $X\in g(\Lambda)$. Similarly, $\Lambda$ has LFC if and only if (3.2) holds for all $X,Y\in g(\Lambda)$. Hence the canonical map $\mathrm{LF}_{1}(\Lambda)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $[X]_{\mathrm{iso}}\mapsto[X]_{\mathrm{st}}$ is surjective by definition of $\mathrm{Cl}(\Lambda)$ and is bijective precisely when $\Lambda$ has LFC. Moreover, $\Lambda$ has SFC precisely when the fibre above 00 has exactly one element, that is, when every stably free $\Lambda$-lattice of rank $1$ is in fact free (of rank $1$).
@@
The claim for LFC was proven by Fröhlich [Frö75, VIII], whose approach was subsequently refined by Swan. Here we adapt the proof of [Nic21a, Theorem 1.1]. By [Swa83, Theorem A10] the diagram
@@
Let $\Lambda\subset\Gamma$ be $R$-orders such that $K\Lambda=K\Gamma$. If $\Lambda$ has SFC (resp. LFC) then $\Gamma$ has SFC (resp. LFC).
@@ 4. Review of cancellation properties for integral group rings
The finite non-cyclic subgroups of $\mathbb{H}^{\times}$ are called the binary polyhedral groups. These are the generalised quaternion groups $Q_{4n}$ for $n\geq 2$ or one of $\tilde{T}$, $\tilde{O}$ or $\tilde{I}$, the binary tetrahedral, octahedral and icosahedral groups (see [CR87, §51A] and the references therein). It is well-known that a finite group $G$ has a binary polyhedral quotient precisely when the group algebra $\mathbb{Q}[G]$ fails the Eichler condition (see [Nic21a, Proposition 1.5] or [CR87, (51.3)], for example).
We now recall two important results of Swan [Swa83, Theorems I and II].
@@
Suppose that $G$ has no binary polyhedral quotient other than those listed in (4.1). We can write $\mathcal{M}=\mathcal{M}_{1}\oplus\cdots\oplus\mathcal{M}_{r}$, where each $\mathcal{M}_{i}$ is a maximal $R$-order contained in a simple $K$-algebra $A_{i}$, and each $A_{i}$ either satisfies the Eichler condition or is a totally definite quaternion algebra. By Corollary 3.3, it suffices to show that each $\mathcal{M}_{i}$ has LFC. Fix $i$ with $1\leq i\leq r$. If $A_{i}$ satisfies the Eichler condition then $\mathcal{M}_{i}$ has LFC by Theorem  3.4. So suppose that $A_{i}$ is a totally definite quaternion algebra. Let $\psi_{i}|_{G}:G\rightarrow\mathbb{H}^{\times}$ denote group homomorphism obtained by restricting the composition $\psi_{i}:\mathbb{Q}[G]\twoheadrightarrow A_{i}\hookrightarrow\mathbb{H}$. Let $H_{i}$ and $N_{i}$ denote the image and kernel of $\psi_{i}|_{G}$, respectively. Then the elements of $H_{i}$ span $\mathbb{H}$ over $\mathbb{R}$ and so $H_{i}$ must be non-abelian since $\mathbb{H}$ is a noncommutative ring. Thus $H_{i}$ is a binary polyhedral quotient of $G$ and so by hypothesis $H_{i}$ is isomorphic to one of the groups listed in (4.1). Let $e_{N_{i}}=|N_{i}|^{-1}\sum_{n\in N_{i}}n$ denote the central idempotent of $\mathbb{Q}[G]$ attached to $N_{i}$. Then
@@
The group ring $\mathbb{Z}[G]$ fails SFC if $G$ is one of the following $5$ groups:
$Q_{12}\times C_{2},\,Q_{16}\times C_{2},\,Q_{20}\times C_{2},\,G_{(36,7)},G_{(100,7)}.$
@@ 5.2. Review of results of Reiner–Ullom
Then the following statements hold.
@@
    Let $\Lambda^{\prime}$ be a maximal $R$-order in $A$ containing $\Lambda$. Then
    $\Lambda^{\prime}\otimes_{\Lambda}M(u)\cong\Lambda^{\prime}\cdot M(u)=\Lambda^{\prime}.$
@@
In the setting of Proposition 5.2, the following statements hold.
@@
Parts (i) and (ii) follow from Proposition 5.2 parts (ii) and (v), respectively. Part (iii) follows from the definition of locally free class group given in §2.5. Part (iv) follows from parts (ii) and (iii). ∎
@@ 5.3. A reduction criterion for SFC
Assume the setting of Corollary 5.3. Suppose that both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and that $W$ is a group. Let $N$ be a normal subgroup of $\ker(\partial)$ contained in $W$ and let $S$ be a generating subset of $\ker(\partial)/N$. Then the following assertions are equivalent.
@@ 5.4. A criterion in terms of sizes of locally free class groups
As shown in the proof of loc. cit. $\partial^{\prime}$ can be defined by $\partial=\partial^{\prime}\circ\pi$ and taking lifts via $\pi$. (Note that the map $\partial$ here is different to that of loc. cit.) Thus $\mathrm{image}(\partial)=\mathrm{image}(\partial^{\prime})$. Therefore the desired result follows from the existence and exactness of (5.4). ∎
@@
where $J(\mathfrak{R})$ denotes the Jacobson radical of $\mathfrak{R}$ (see [CR81, Exercise 5.2], for example). In particular, if $\mathfrak{R}$ is finite then |×$|\mathfrak{R}^{\times}|=|J(\mathfrak{R})||(\mathfrak{R}/J(\mathfrak{R}))^{\times}|$. If $\mathfrak{R}$ is a finite-dimensional algebra over a finite field, then $J(\mathfrak{R})$ can be computed using the algorithm of [FR85].

### v2

@@ Abstract.
Let $K$ be a number field, let $A$ be a finite-dimensional semisimple $K$-algebra, and let $\Lambda$ be an $\mathcal{O}_{K}$-order in $A$. We give practical algorithms that determine whether or not $\Lambda$ has stably free cancellation (SFC). As an application, we determine all finite groups $G$ of order at most $383$ such that the integral group ring $\mathbb{Z}[G]$ has SFC.
@@ 1. Introduction
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$. Let $A$ be a finite-dimensional semisimple $K$-algebra and let $\Lambda$ be an $\mathcal{O}$-order in $A$. For example, if $G$ is a finite group then the group ring $\mathcal{O}[G]$ is an $\mathcal{O}$-order in the group algebra $K[G]$. For $k\geq 0$, let $\Lambda^{(k)}$ denote the (left) $\Lambda$-module that is the direct sum of $k$ copies of $\Lambda$. The order $\Lambda$ is said to have stably free cancellation (SFC) if every finitely generated stably free $\Lambda$-module is in fact free, that is, if $X$ is a finitely generated $\Lambda$-module such that $X\oplus\Lambda^{(n)}\cong\Lambda^{(n+m)}$ for some $m,n\geq 0$, then $X\cong\Lambda^{(m)}$.
@@
In this article, we show that the question of whether an arbitrary order $\Lambda$ has SFC or not can in principle always be decided algorithmically. In fact, we present three algorithms, all of which are practical in certain situations. The first of these, Algorithm 8.9, is only practical when $A$ is of ‘small’ dimension. The second, Algorithm 9.1, cannot prove that $\Lambda$ has SFC, but can be used to show that it fails SFC, and in this situation is often much faster than Algorithm 8.9. The third, Algorithm 10.3, uses fibre products to reduce the problem for the original order to that for an order in an algebra of smaller dimension as well as the computation of a certain subgroup of the unit group of a finite ring. These algorithms require several other algorithms of independent interest, including those that can be used to compute certain primary decompositions and unit groups of finite rings (see §11), and new freeness tests for certain modules over orders (see §12).
Orders of particular interest are integral group rings $\mathbb{Z}[G]$ for $G$ a finite group. For instance, the property of whether or not $\mathbb{Z}[G]$ has SFC or LFC has applications in topology (see [Joh03], [Nic21b], [Nic24a], for example) and in the determination of whether a finite Galois extension of $\mathbb{Q}$ has a normal integral basis or not (see [Cou94], [Cou98]).
@@
A special case of a result due to Swan and Fröhlich says that if $\mathbb{Z}[G]$ has SFC then $\mathbb{Z}[G/N]$ has SFC for every quotient $G/N$, or equivalently, if $\mathbb{Z}[G/N]$ fails SFC for some choice of $N$, then $\mathbb{Z}[G]$ also fails SFC (see Theorem 3.1 and Lemma 4.1). The following result is Theorem 9.4, in which the claim that each $\mathbb{Z}[G]$ fails SFC was proven using Algorithm 9.1; the claim that $\mathbb{Z}[G/N]$ has SFC for each proper quotient $G/N$ can be interpreted as saying that each $G$ is minimal with respect to quotients.
@@
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}^{2}$, or
@@
$\tilde{T}\times C_{2},\,\tilde{T}\times Q_{12},\,\tilde{T}\times Q_{20},\,Q_{8}\rtimes\tilde{T},\,Q_{8}\rtimes Q_{12},\,\tilde{I}\times C_{2},\,G_{(192,183)},\,G_{(384,580)},\,G_{(480,962)}.$
@@
    $\tilde{T}\times C_{2}^{2}$, $\tilde{O}\times C_{2}$, $\tilde{I}\times C_{2}^{2}$,
@@
This represents a significant advance in the state-of-the-art, since before the results of this article, such a classification was only known for groups $G$ with $|G|\leq 31$ (for example, it was not known whether $\mathbb{Z}[G_{(32,14)}]$ had SFC or not). In fact, Theorem 1.4 is a special case of Theorem 13.4, which gives a classification for $|G|\leq 1023$, but with $49$ explicitly listed gaps. In other words, of the $11759892$ groups $G$ with $|G|\leq 1023$, there are now only $49$ for which it is not known whether $\mathbb{Z}[G]$ has SFC or not.
Nicholson [Nic24b, Theorem 1.7] has now shown that $\mathbb{Z}[\tilde{T}\times C_{2}]$ in fact has LFC, using our result that it has SFC (i.e. part of Theorem 1.3) as an input. By combining this with further new results on integral group rings with LFC, and using Theorem 1.1 as a crucial input, he has given classifications of all integral group rings $\mathbb{Z}[G]$ with LFC and SFC, provided that either (i) $G$ has no quotient isomorphic to $\tilde{T}$, $\tilde{O}$ or $\tilde{I}$, or (ii) $G$ does have a quotient isomorphic to $C_{2}\times C_{2}$ (see [Nic24b, Theorems A and B]). We remark that in the case of integral group rings, the LFC property discussed here is equivalent to the projective cancellation (PC) property discussed in loc. cit.
@@ Acknowledgements
The authors are grateful to John Nicholson for helpful comments and discussions, including suggestions for which integral group rings to investigate initially. They would also like to thank Nigel Byott for several corrections and suggestions. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any author accepted manuscript version arising.
@@ 2.5. Locally free class groups
where $X_{1},X_{2},X_{3}\in g(\Lambda)$. This is a well-defined commutative and associative operation, with identity element $[\Lambda]_{\mathrm{st}}$. Moreover, inverses exist since by (2.1), given $X\in g(\Lambda)$ there exists $X^{\prime}\in g(\Lambda)$ such that $X\oplus X^{\prime}\cong\Lambda^{(2)}$. Therefore $\mathrm{Cl}(\Lambda)$ is an abelian group. Moreover, $\mathrm{Cl}(\Lambda)$ is finite since $\mathrm{LF}_{1}(\Lambda)$ is finite and the canonical map $\mathrm{LF}_{1}(\Lambda)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $[X]_{\mathrm{iso}}\mapsto[X]_{\mathrm{st}}$ is surjective.
@@ 3.1. Stably free and locally free cancellation
By (2.2), $\Lambda$ has SFC if and only if (3.1) holds for all $X\in g(\Lambda)$. Similarly, $\Lambda$ has LFC if and only if (3.2) holds for all $X,Y\in g(\Lambda)$. Hence the canonical surjective map $\mathrm{LF}_{1}(\Lambda)\rightarrow\mathrm{Cl}(\Lambda)$ defined by $[X]_{\mathrm{iso}}\mapsto[X]_{\mathrm{st}}$ is bijective precisely when $\Lambda$ has LFC. Moreover, $\Lambda$ has SFC precisely when the fibre above 00 has exactly one element, that is, when every stably free $\Lambda$-lattice of rank $1$ is in fact free (of rank $1$).
@@
The claim for LFC was proven by Fröhlich [Frö75, VIII], whose approach was subsequently refined by Swan. Here we adapt the proof of [Nic21a, Theorem 1.1], which shows the result for SFC in the case that $\Lambda$ and $\Gamma$ are integral group rings. By [Swa83, Theorem A10] the diagram
@@
Let $\Lambda\subseteq\Gamma$ be $R$-orders such that $K\Lambda=K\Gamma$. If $\Lambda$ has SFC (resp. LFC) then $\Gamma$ has SFC (resp. LFC).
@@ 4. Review of cancellation properties for integral group rings
The finite non-cyclic subgroups of $\mathbb{H}^{\times}$ are called the binary polyhedral groups. These are the generalised quaternion groups $Q_{4n}$ for $n\geq 2$ or one of $\tilde{T}$, $\tilde{O}$ or $\tilde{I}$, the binary tetrahedral, octahedral and icosahedral groups (see [CR87, §51A] and the references therein). It is well known that a finite group $G$ has a binary polyhedral quotient precisely when the group algebra $\mathbb{Q}[G]$ does not satisfy the Eichler condition (see [Nic21a, Proposition 1.5] or [CR87, (51.3)], for example).
We now recall two important theorems of Swan [Swa83, Theorems I and II].
@@
Corollary 4.3 (Swan).
Let $G$ be a finite group with a quotient isomorphic to $Q_{4n}$ for some $n\geq 6$. Then $\mathbb{Z}[G]$ fails SFC.
Proof.
This follows from one direction of Theorem 4.2 combined with Lemma 4.1. ∎
@@
Remark 4.5.
In order, the groups listed in (4.1) are isomorphic to
$G_{(8,4)},\,G_{(12,1)},\,G_{(16,9)},\,G_{(20,1)},\,G_{(24,4)},\,G_{(28,1)},\,G_{(36,1)},\,G_{(60,3)},$
$G_{(24,3)}\cong\mathrm{SL}_{2}(\mathbb{F}_{3}),\quad G_{(48,28)}\cong\mathrm{CSU}_{2}(\mathbb{F}_{3}),\quad G_{(120,5)}\cong\mathrm{SL}_{2}(\mathbb{F}_{5}).$
@@
Suppose that $G$ has no binary polyhedral quotient other than those listed in (4.1). We can write $\mathcal{M}=\mathcal{M}_{1}\oplus\cdots\oplus\mathcal{M}_{r}$, where each $\mathcal{M}_{i}$ is a maximal $R$-order contained in a simple $K$-algebra $A_{i}$, and each $A_{i}$ either satisfies the Eichler condition or is a totally definite quaternion algebra. By Corollary 3.3, it suffices to show that each $\mathcal{M}_{i}$ has LFC. Fix $i$ with $1\leq i\leq r$. If $A_{i}$ satisfies the Eichler condition then $\mathcal{M}_{i}$ has LFC by Theorem  3.4. So suppose that $A_{i}$ is a totally definite quaternion algebra. Let $\psi_{i}|_{G}:G\rightarrow\mathbb{H}^{\times}$ denote the group homomorphism obtained by restricting the composition $\psi_{i}:\mathbb{Q}[G]\twoheadrightarrow A_{i}\hookrightarrow\mathbb{H}$. Let $H_{i}$ and $N_{i}$ denote the image and kernel of $\psi_{i}|_{G}$, respectively. Then the elements of $H_{i}$ span $\mathbb{H}$ over $\mathbb{R}$ and so $H_{i}$ must be non-abelian since $\mathbb{H}$ is a noncommutative ring. Thus $H_{i}$ is a binary polyhedral quotient of $G$ and so by hypothesis $H_{i}$ is isomorphic to one of the groups listed in (4.1). Let $e_{N_{i}}=|N_{i}|^{-1}\sum_{n\in N_{i}}n$ denote the central idempotent of $\mathbb{Q}[G]$ attached to $N_{i}$. Then
@@
Let $G$ be one of the following $5$ groups:
$Q_{12}\times C_{2},\quad Q_{16}\times C_{2},\quad Q_{20}\times C_{2},\quad G_{(36,7)},\quad G_{(100,7)}.$
Then $\mathbb{Z}[G]$ fails SFC.
@@ 5.2. Review of results of Reiner–Ullom
Then the following statements hold:
@@
    Let $\mathcal{M}$ be a maximal $R$-order in $A$ containing $\Lambda$. Then
    $\mathcal{M}\otimes_{\Lambda}M(u)\cong\mathcal{M}\cdot M(u)=\mathcal{M}.$
@@
In the setting of Proposition 5.2, the following statements hold:
@@
Parts (i) and (ii) follow from Proposition 5.2 parts (ii) and (v), respectively. Part (iii) follows from the definition of $\mathrm{Cl}(\Lambda)$ given in §2.5. Part (iv) follows from parts (ii) and (iii). ∎
@@ 5.3. A reduction criterion for SFC
Assume the setting of Corollary 5.3. Suppose that both $\Lambda_{1}$ and $\Lambda_{2}$ have SFC and that $W$ is a group. Let $N$ be a normal subgroup of $\ker(\partial)$ contained in $W$ and let $S$ be a generating subset of $\ker(\partial)/N$. Then the following assertions are equivalent:
@@ 5.4. A criterion in terms of sizes of locally free class groups
As shown in the proof of loc. cit. $\partial^{\prime}$ can be defined by $\partial=\partial^{\prime}\circ\pi$ and taking lifts via $\pi$. (Note that the map $\partial$ here is different to that of loc. cit.) Thus $\mathrm{im}(\partial)=\mathrm{im}(\partial^{\prime})$. Therefore the desired result follows from the existence and exactness of (5.4). ∎
@@
where $J(\mathfrak{R})$ denotes the Jacobson radical of $\mathfrak{R}$ (see [CR81, Exercise 5.2], for example). In particular, if $\mathfrak{R}$ is finite then |×$|\mathfrak{R}^{\times}|=|J(\mathfrak{R})||(\mathfrak{R}/J(\mathfrak{R}))^{\times}|$. If $\mathfrak{R}$ is a finite-dimensional algebra over a finite field, then $J(\mathfrak{R})$ can be computed using the algorithm of [FR85]. We will use these observations for the computation of the cardinality of $\overline{\Lambda}^{\times}$ in certain situations.